OpenStudy (anonymous):

a chemist neds 90 milliliters of a 71% solution but has only 67% and 76% solutions available. Find how many milliliters of each that should be mixed to get the desired solution a) 37 ml of 67%; 53 ml of 76% b) 50ml of 67%; 40 ml of 76% c) 40 ml of 67%; 50 ml of 76% d) 53 ml of 67%; 37 ml of 76% i need process please!

OpenStudy (anonymous):

put \(x\) as the number of ml of 76% solution so \(90-x\) is the number of ml of 67% solution you want the total to be 71% of 90 \(.71\times 90=63.9\) so solve \[.76x+.67(90-x)=63.9\] and solve for \(x\)

OpenStudy (anonymous):

i dont know how to please explain to mee!

OpenStudy (anonymous):

how to solve for \(x\)? probably easiest to multiply both sides by 100 to get rid of annoying decimals and write \[.76x+.67(90-x)=63.9\] \[76x+67(90-x)=6390\] multiply out and get \[76x+6030-67x=6390\] combine like terms and get \[9x+6030=6390\] subtract 6030 from both sides and get \[9x=360\] divide by 9 and get \[x=40\]

OpenStudy (anonymous):

now don't forget that line number one i wrote was put \(x \) as the number of ml of 76% so that means 40 ml of the 76% solution and the other 50 ml of the 67% solution

OpenStudy (anonymous):

thank you so muchhh <3

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

please can you help me in an other :$